OpenStudy (barrelracer17):

Select interior, exterior, or on the circle (x - 5) 2 + (y + 3) 2 = 25 for the following point. (2, 3)

OpenStudy (misty1212):

HI!

OpenStudy (misty1212):

do you know the center and radius of this circle (hint, it is in the right form for you to eyeball it )

OpenStudy (barrelracer17):

is it 25?

OpenStudy (misty1212):

is what 25?

OpenStudy (barrelracer17):

the radius.. i dont know im not so good with geometry :/

OpenStudy (misty1212):

general form of a circle with center \((h,k)\) and radius \(r\) is \[(x-h)^2+(y-k)^2=r^2\]

OpenStudy (misty1212):

you have \[(x-5)^2+(y+3)^2=25\] so you can read off the center and radius from the equation

OpenStudy (barrelracer17):

so is it interior?

OpenStudy (barrelracer17):

Or exterior?